# Lite Any Stereo: Efficient Zero-Shot Stereo Matching

### v1

@@ 1 Introduction
From the foundational work of [marr1988cooperative] through the classical advances culminating in works such as [taniai2017continuous], stereo vision has seen decades of steady progress built on a wide range of algorithmic ideas. In the past decade, deep learning methods have driven a dramatic leap in accuracy—leading many researchers to view stereo vision as largely solved—yet these gains often rely on large, computationally expensive models that are difficult to deploy on resource-limited hardware.
Learning-based methods [lipson2021raft, li2022practical, xu2023accurate, wang2024selective] achieve remarkable accuracy, continuously raising scores on standard benchmarks [middlebury, eth3d, kitti12, kitti15]. These methods can be broadly categorized as accuracy-oriented. More recently, the emergence of foundation models trained on internet-scale data, such as the DepthAnything series [yang2024depth, yang2024depth2], has led to further advancement in the field. Stereo models augmented with depth priors [wen2025foundationstereozeroshotstereomatching, cheng2025monstermarrymonodepthstereo, jiang2025defomstereodepthfoundationmodel] have demonstrated excellent zero-shot generalization: a single set of model weights can deliver strong results across diverse scenarios. However, despite these successes, such approaches prioritize accuracy over efficiency, which limits their practical applications.
In contrast, efficiency-oriented approaches [shamsafar2022mobilestereonet, guo2024lightstereochannelboostneed, xu2025banet] trade accuracy for faster inference and lower resource use, however, the accuracy gap to large models remains significant. The absence of light models with strong zero-shot generalization may lead to a conclusion that such models lack capacity for zero-shot applications. Consequently, most approaches still rely on domain-specific fine-tuning, thus falling short of being practical off-the-shelf solutions.
Training from large data is a key factor in enabling zero-shot capability. StereoAnything [guo2024stereo] scaled training by using a monocular depth model [yang2024depth2] to generate pseudo-disparity maps, yielding 30M real-world samples. While large data improves zero-shot performance for efficient models [guo2024lightstereochannelboostneed], they remain far behind accuracy-oriented models, primarily due to the limited quality of monocular depth.
@@ 2.1 Accurate Stereo Methods
Modern stereo matching methods based on deep learning typically construct either 3D or 4D cost volumes to capture pixel-level correspondences between the left and right views. These volumes serve as the foundation for disparity estimation, with convolutional neural networks used to aggregate contextual information. Broadly, CNN-based approaches can be categorized into two types: those that emphasize cost aggregation [pang2017cascade, chang2018pyramid, liang2018learning, yang2019hierarchical, zhang2019ga, guo2019group, xu2020aanet, cheng2020hierarchical, gu2020cascade, tankovich2021hitnet, guo2023openstereo], and those that adopt iterative refinement strategies [lipson2021raft, li2022practical, xu2023iterative, wang2024selective] to progressively enhance disparity predictions. Another line of research explores transformer-based stereo models [guo2022context, li2021revisiting, su2022chitransformer, weinzaepfel2023croco, xu2023unifying], which leverage attention mechanisms to capture global context. In addition, some methods explicitly address temporal consistency in stereo videos [karaev2023dynamicstereo, jing2024matchstereovideos, jing2024matchstereovideosbidirectional, jing2025stereovideotemporallyconsistent], aiming to produce consistent disparities across time. While these models perform well on domain-specific benchmarks [kitti12, kitti15, middlebury, eth3d], generalization to unseen domains in zero-shot settings remains a major challenge. To address this, some works have focused on learning domain-invariant representations [chuah2022itsa, zhang2022revisiting, chang2023domain, rao2023masked] and robust modules [shen2021cfnet, Jing_2023_ICCV, guo2024stereo], aiming to improve performance under domain shifts. A recent new direction involves leveraging prior knowledge from monocular foundation models [yang2024depth, yang2024depth2]. By integrating monocular depth cues into stereo pipelines, several methods [bartolomei2024stereo, zhang2024learning, zhou2024all, cheng2025monstermarrymonodepthstereo, jiang2025defomstereodepthfoundationmodel, wen2025foundationstereozeroshotstereomatching] demonstrate enhanced zero-shot generalization. However, these improvements often come at the cost of increased computational complexity, limiting their deployment in real-time or resource-constrained scenarios.
@@ 2.2 Efficient Stereo Matching
Real-time performance is critical for stereo matching in practical applications. Early approaches [khamis2018stereonet, duggal2019deeppruner, wang2020fadnet] address this by estimating disparities at reduced spatial resolutions to lower computational cost, though at the expense of accuracy. Subsequent methods aim to improve this efficiency and accuracy trade-off based on 3D aggregation. Representative advancements include guided cost volume excitation in CoEx [bangunharcana2021correlate], edge-aware upsampling in BGNet [xu2021bilateral], and sparse attention for selective high-resolution matching in Fast-ACVNet [xu2022acvnet, xu2023accurate]. To further reduce complexity, another line of works have explored 2D-based alternatives to traditional 3D convolutions. Notable examples include AANet [xu2020aanet], which employs deformable 2D convolutions for adaptive cost aggregation; HITNet [tankovich2021hitnet], which reconstructs disparities through iterative warping without explicit cost volumes; and MobileStereoNet-2D [shamsafar2022mobilestereonet], which adopts lightweight MobileNet blocks. More recent efforts continue this trend with specialized strategies: uncertainty-aware adaptive warping in Lite-CREStereo++ [Jing_2023_ICCV], channel-wise enhancement in LightStereo [guo2024lightstereochannelboostneed], and frequency-disparity guided bilateral aggregation in BANet [xu2025banet]. Despite these advances in efficiency, most of these models are optimized for a specific domain, particularly the KITTI online benchmarks [kitti12, kitti15], and still struggle to generalize well across diverse, unseen scenarios. Achieving strong zero-shot ability while maintaining real-time efficiency remains an open challenge.
@@ 3.1 Framework
Feature Extraction. Recent methods [jiang2025defomstereodepthfoundationmodel, cheng2025monstermarrymonodepthstereo, wen2025foundationstereozeroshotstereomatching] have demonstrated remarkable improvements by leveraging depth features from DepthAnything (DA) [yang2024depth, yang2024depth2]. However, we find that the additional computational overhead introduced by DA, even with the smallest DA-S variant, is prohibitive for an efficiency-oriented stereo model. Therefore, we adopt a conventional backbone [guo2024lightstereochannelboostneed, xu2025banet] pretrained on ImageNet [deng2009imagenet] without external priors for feature extraction. Moreover, we find that compared with more recent CNN backbones [liu2022convnet, woo2023convnext], the channel configuration of MobileNetV2 [sandler2018mobilenetv2] is better aligned with the requirements of our task. Specifically, given a pair of rectified stereo images $\{\mathbf{I}_{L},\mathbf{I}_{R}\}\in\mathbb{R}^{H\times W\times 3}$, we employ two weight-sharing networks to pyramidally generate multi-scale features $\{\mathbf{F}_{L}^{s}\}$ and $\{\mathbf{F}_{R}^{s}\}$, where $s\in\left\{\frac{1}{4},\frac{1}{8},\frac{1}{16},\frac{1}{32}\right\}$ denotes the downsampling ratio. To unify the spatial resolution for subsequent processing, features at all scales are upsampled to $\tfrac{1}{4}$ resolution using residual upsampling blocks, following [guo2024lightstereochannelboostneed].
@@
Cost Aggregation. In recent years, 2D cost aggregation has replaced 3D cost aggregation in efficient stereo models [shamsafar2022mobilestereonet, guo2024lightstereochannelboostneed, xu2025banet]. While this reduces computational cost, the limited receptive field of 2D convolutions in the spatial dimension restricts model’s zero-shot ability. Towards this, we introduce a hybrid aggregation module that complements 2D and 3D representations: 3D block $\mathbf{G}_{\text{3D}}(\cdot)$ provides disparity-dimension perception, while 2D block $\mathbf{G}_{\text{2D}}(\cdot)$ enhances spatial-dimension perception.
Specifically, multi-scale 3D convolutions are employed in the 3D block, and ConvNeXt layers [liu2022convnet] are used in the 2D block. A naive half–half hybrid design is ineffective, as 3D convolutions dominate the compute budget along the disparity dimension, even though many disparity levels contribute little to the final results. Instead, we retain only a small proportion of 3D component to preserve disparity perception while maintaining efficiency. We also explore multiple design choices for integrating 2D and 3D branches while keeping the overall computational cost fixed, as illustrated in Fig. 5. In particular: (a) follows [xu2025banet], where bilateral aggregation is adopted and the outputs of 2D and 3D branches are summed; (b) and (c) apply serial connections, but with different ordering of 2D and 3D blocks; (d) interleaves 2D and 3D blocks. Through ablation studies (Section 4.3), we find that design (c) achieves the best performance. Therefore, we adopt (c) as the default configuration of our cost aggregation module, formulated as follows,
@@
Disparity Estimation. Similar to other efficient methods [guo2024lightstereochannelboostneed, xu2025banet], we apply the soft-argmax operation to regress the disparity map $\mathbf{d}$ at $\frac{1}{4}$ scale:
@@
Flickr1024 [Flickr1024] 0.73 1K
InStereo2k [bao2020instereo2k] 0.93 2K
Holopix50K [hua2020holopix50k] 0.74 49K
Driving Stereo [yang2019drivingstereo] 0.40 174K
SouthKenSV [jing2024matchstereovideosbidirectional] 0.92 113K
UASOL [bauer2019uasol] 2.74 156K
@@
Table 2: Ablation study on KITTI 2012 [kitti12] (K.12), KITTI 2015 [kitti15] (K.15), ETH3D [eth3d] (E.), and Middlebury [middlebury] (M.). We report D1 for KITTI, Bad 1 for ETH3D, and Bad 2 for Middlebury. We also report MACs (Multiply-Accumulate Operations) for backbone design choices. Unless otherwise specified, models are trained for 200K iterations without data augmentation on a 1.4M-image subset from synthetic datasets [wen2025foundationstereozeroshotstereomatching, li2022practical, cabon2020vkitti2, fallingthings, mayer2016large] using the default operations in [xu2025banet]. Default settings in final model are marked in gray.
@@ 3.2 Training Strategy
To achieve zero-shot capability, the proposed model is trained on a large amount of high-quality data by following carefully designed training stages, which we illustrate in Fig. 4. We compile datasets into two categories: synthetic annotated data and realistic unannotated data. Since synthetic datasets provide accurate ground-truth annotations, we first train our model on them from scratch to enhance model’s matching ability. Our synthetic datasets combine SceneFlow [mayer2016large] (35K), FallingThings [fallingthings] (30K), FSD [wen2025foundationstereozeroshotstereomatching] (1.1M), CREStereo [li2022practical] (0.2M), VKITTI2 [cabon2020vkitti2] (21K), TartanAir [tartanair2020iros] (0.31M), and Dynamic Replica [karaev2023dynamicstereo] (0.14M), resulting in a total of 1.8M annotated samples. While other synthetic datasets are available, such as IRS [wang2019irs], Sintel [sintel], Spring [Mehl2023_Spring], and InfinigenSV [jing2024matchstereovideosbidirectional], they are excluded due to annotation quality issues (e.g., inaccurate disparities for transparent objects in IRS) or domain mismatch (e.g., cinematic or natural scenes in Spring and InfinigenSV).
@@
where $F_{i}$ and $F^{\prime}_{i}$ are feature vectors from the teacher and student models, respectively. Here, we evaluate several distillation schemes: (a) training only the student model with fixed teacher model weights, (b) updating the teacher model via Exponential Moving Average (EMA) [polyak1992acceleration], and (c) directly copying student model weights to the teacher model at each iteration. Our ablation study (Section 4.3) shows that the simplest strategy (a) yields the best performance, and is therefore adopted in the second stage.
High-quality real-world stereo data with annotations remain scarce and sparse (e.g., Lidar-based ground truth), which constrains training scalability. In contrast, the vast amount of unlabeled real-world stereo data has not been used by the community. With the increasing availability and diversity of such data, in Stage ③, we leverage unannotated stereo pairs to improve model generalization, collecting a total of 0.5M pairs, as summarized in Tab. 1. Pseudo labels are generated using FoundationStereo [wen2025foundationstereozeroshotstereomatching], a strong accurate model. Importantly, even for datasets with ground-truth annotations (e.g., DrivingStereo [yang2019drivingstereo]), we use pseudo but dense labels for stereo pairs (excluding the weather subset), as they provide stronger supervision than sparse data.
We further observe that data quality is more critical than scale: incorporating low-quality or domain-specific data in the third stage can degrade zero-shot performance. For example, Stereo4D [jin2024stereo4d] contains 18M stereo pairs mined from internet videos but only at a limited resolution ($\times512$), HRWSI [hrwsi] suffers from poor rectification quality, and datasets such as SCOD [scod] are restricted to narrow domains. Finally, we do not apply self-distillation here, as it provides no observable performance gains when training with pseudo labels. As shown in Fig. 6, the proposed strategy progressively improves the disparity estimation, making the leaves and background regions noticeably clearer.
@@ 4 Experiments
Table 3: Zero-shot generalization results on four public benchmarks: KITTI 2012 [kitti12], KITTI 2015 [kitti15], ETH3D [eth3d], and Middlebury (H) [middlebury]. The most commonly used metrics are adopted. In the first block, all efficient methods are trained only on Scene Flow [mayer2016large]. In the second and third blocks, methods are allowed to train on any existing datasets excluding the four target domains. Accuracy-based methods are shown as reference. The weights and parameters are fixed for evaluation. MACs (G) are measured at the KITTI resolution of $1242\times 375$. indicates models retrained with original code on the same synthetic data as ours. denotes results trained on 30M pseudo-labeled samples using the strategy in [guo2024stereo]. marks results reported by [wen2025foundationstereozeroshotstereomatching]. The best and second best are marked with colors.
@@
CoEX [bangunharcana2021correlate] 22.30 3.37 17.33 3.00 31.97 14.05 26.42 4.90 54
MobileStereoNet-2D [shamsafar2022mobilestereonet] 19.30 2.51 21.88 2.85 17.10 1.83 37.98 7.54 127
MobileStereoNet-3D [shamsafar2022mobilestereonet] 19.59 2.67 17.89 3.01 18.42 3.02 25.26 4.61 564
FastACV [xu2022acvnet] 13.90 2.05 11.83 2.21 7.84 1.05 19.61 4.66 72
FastACV+ [xu2023accurate] 16.69 2.29 15.28 3.27 11.50 2.43 27.34 7.16 85
Lite-CREStereo++ [Jing_2023_ICCV] 5.93 1.29 7.37 1.36 8.95 1.18 14.91 3.32 101
LightStereo-M [guo2024lightstereochannelboostneed] 6.76 1.26 6.79 1.42 13.93 0.73 16.99 2.06 33
LightStereo-L [guo2024lightstereochannelboostneed] 6.80 1.27 6.62 1.29 9.66 0.50 17.23 2.88 84
BANet-2D [xu2025banet] 14.75 2.23 16.98 3.91 44.89 35.95 26.79 6.96 36
BANet-3D [xu2025banet] 17.39 2.41 17.28 3.36 29.27 14.06 28.79 8.05 78
@@
LightStereo-M [guo2024lightstereochannelboostneed] 4.10 0.99 4.97 1.13 5.33 0.41 10.85 1.51 33
BANet-2D [xu2025banet] 3.90 0.93 4.71 1.07 5.92 0.38 10.05 1.34 36
StereoAnything-L [guo2024stereo] 4.00 0.92 4.81 1.10 3.81 0.31 9.82 1.21 84
@@
Selective-IGEV [wang2024selective] 3.20 4.50 3.40 7.50 3619
FoundationStereo [wen2025foundationstereozeroshotstereomatching] 2.51 0.67 2.83 0.86 0.49 0.14 1.12 0.37 12824
Table 4: Comparison of model performance on DrivingStereo weather [yang2019drivingstereo]. Lower value is better for both metrics.

### v2

@@ 1 Introduction
From the foundational work of [37] through the classical advances culminating in works such as [51], stereo vision has seen decades of steady progress built on a wide range of algorithmic ideas. In the past decade, deep learning methods have driven a dramatic leap in accuracy—leading many researchers to view stereo vision as largely solved—yet these gains often rely on large, computationally expensive models that are difficult to deploy on resource-limited hardware.
Learning-based methods [35, 32, 68, 58] achieve remarkable accuracy, continuously raising scores on standard benchmarks [46, 47, 14, 40]. These methods can be broadly categorized as accuracy-oriented. More recently, the emergence of foundation models trained on internet-scale data, such as the DepthAnything series [73, 74], has led to further advancement in the field. Stereo models augmented with depth priors [61, 9, 23] have demonstrated excellent zero-shot generalization: a single set of model weights can deliver strong results across diverse scenarios. However, despite these successes, such approaches prioritize accuracy over efficiency, which limits their practical applications.
In contrast, efficiency-oriented approaches [48, 20, 66] trade accuracy for faster inference and lower resource use, however, the accuracy gap to large models remains significant. The absence of light models with strong zero-shot generalization may lead to a conclusion that such models lack capacity for zero-shot applications. Consequently, most approaches still rely on domain-specific fine-tuning, thus falling short of being practical off-the-shelf solutions.
Training from large data is a key factor in enabling zero-shot capability. StereoAnything [19] scaled training by using a monocular depth model [74] to generate pseudo-disparity maps, yielding 30M real-world samples. While large data improves zero-shot performance for efficient models [20], they remain far behind accuracy-oriented models, primarily due to the limited quality of monocular depth.
@@ 2.1 Accurate Stereo Methods
Modern stereo matching methods based on deep learning typically construct either 3D or 4D cost volumes to capture pixel-level correspondences between the left and right views. These volumes serve as the foundation for disparity estimation, with convolutional neural networks used to aggregate contextual information. Broadly, CNN-based approaches can be categorized into two types: those that emphasize cost aggregation [42, 7, 34, 71, 76, 21, 70, 10, 15, 52, 17, 18], and those that adopt iterative refinement strategies [35, 32, 67, 58] to progressively enhance disparity predictions. Another line of research explores transformer-based stereo models [16, 33, 50, 60, 69], which leverage attention mechanisms to capture global context. In addition, some methods explicitly address temporal consistency in stereo videos [29, 27, 28, 26], aiming to produce consistent disparities across time. While these models perform well on domain-specific benchmarks [14, 40, 46, 47], generalization to unseen domains in zero-shot settings remains a major challenge. To address this, some works have focused on learning domain-invariant representations [11, 77, 8, 44] and robust modules [49, 25, 19], aiming to improve performance under domain shifts. A recent new direction involves leveraging prior knowledge from monocular foundation models [73, 74]. By integrating monocular depth cues into stereo pipelines, several methods [3, 78, 79, 9, 23, 61] demonstrate enhanced zero-shot generalization. However, these improvements often come at the cost of increased computational complexity, limiting their deployment in real-time or resource-constrained scenarios.
@@ 2.2 Efficient Stereo Matching
Real-time performance is critical for stereo matching in practical applications. Early approaches [30, 13, 55] address this by estimating disparities at reduced spatial resolutions to lower computational cost, though at the expense of accuracy. Subsequent methods aim to improve this efficiency and accuracy trade-off based on 3D aggregation. Representative advancements include guided cost volume excitation in CoEx [1], edge-aware upsampling in BGNet [64], and sparse attention for selective high-resolution matching in Fast-ACVNet [65, 68]. To further reduce complexity, another line of works have explored 2D-based alternatives to traditional 3D convolutions. Notable examples include AANet [70], which employs deformable 2D convolutions for adaptive cost aggregation; HITNet [52], which reconstructs disparities through iterative warping without explicit cost volumes; and MobileStereoNet-2D [48], which adopts lightweight MobileNet blocks. More recent efforts continue this trend with specialized strategies: uncertainty-aware adaptive warping in Lite-CREStereo++ [25], channel-wise enhancement in LightStereo [20], and frequency-disparity guided bilateral aggregation in BANet [66]. Despite these advances in efficiency, most of these models are optimized for a specific domain, particularly the KITTI online benchmarks [14, 40], and still struggle to generalize well across diverse, unseen scenarios. Achieving strong zero-shot ability while maintaining real-time efficiency remains an open challenge.
@@ 3.1 Framework
Feature Extraction. Recent methods [23, 9, 61] have demonstrated remarkable improvements by leveraging depth features from DepthAnything (DA) [73, 74]. However, we find that the additional computational overhead introduced by DA, even with the smallest DA-S variant, is prohibitive for an efficiency-oriented stereo model. Therefore, we adopt a conventional backbone [20, 66] pretrained on ImageNet [12] without external priors for feature extraction. Moreover, we find that compared with more recent CNN backbones [36, 62], the channel configuration of MobileNetV2 [45] is better aligned with the requirements of our task. Specifically, given a pair of rectified stereo images $\{\mathbf{I}_{L},\mathbf{I}_{R}\}\in\mathbb{R}^{H\times W\times 3}$, we employ two weight-sharing networks to pyramidally generate multi-scale features $\{\mathbf{F}_{L}^{s}\}$ and $\{\mathbf{F}_{R}^{s}\}$, where $s\in\left\{\frac{1}{4},\frac{1}{8},\frac{1}{16},\frac{1}{32}\right\}$ denotes the downsampling ratio. To unify the spatial resolution for subsequent processing, features at all scales are upsampled to $\tfrac{1}{4}$ resolution using residual upsampling blocks, following [20].
@@
Cost Aggregation. Recent efficient models adopt 2D cost aggregation [48, 20, 66], which reduces computation by collapsing the disparity into channels and performing spatial-only convolutions. However, this strategy does not explicitly model structured continuity along the disparity dimension, potentially limiting generalization. To address this, we introduce a hybrid aggregation module that combines complementary 2D and 3D reasoning. The 3D block $\mathbf{G}_{\text{3D}}(\cdot)$ jointly aggregates over disparity and spatial dimensions to capture cross-disparity structure, while the 2D block $\mathbf{G}_{\text{2D}}(\cdot)$ provides efficient spatial refinement. This design preserves geometric awareness with minimal overhead.
Specifically, multi-scale 3D convolutions are employed in the 3D block, and ConvNeXt layers [36] are used in the 2D block. A naive half–half hybrid design is ineffective, as 3D convolutions dominate the compute budget along the disparity dimension, even though many disparity levels contribute little to the final results. Instead, we retain only a small proportion of 3D component to preserve disparity perception while maintaining efficiency. We also explore multiple design choices for integrating 2D and 3D branches while keeping the overall computational cost fixed, as illustrated in Fig. 5. In particular: (a) follows [66], where bilateral aggregation is adopted and the outputs of 2D and 3D branches are summed; (b) and (c) apply serial connections, but with different ordering of 2D and 3D blocks; (d) interleaves 2D and 3D blocks. Through ablation studies (Section 4.3), we find that design (c) achieves the best performance. Therefore, we adopt (c) as the default configuration of our cost aggregation module, formulated as follows,
@@
Disparity Estimation. Similar to other efficient methods [20, 66], we apply the soft-argmax operation to regress the disparity map $\mathbf{d}$ at $\frac{1}{4}$ scale:
@@
Flickr1024 [59] 0.73 1K
InStereo2k [2] 0.93 2K
Holopix50K [22] 0.74 49K
Driving Stereo [72] 0.40 174K
SouthKenSV [28] 0.92 113K
UASOL [4] 2.74 156K
@@
Table 2: Ablation study on KITTI 2012 [14] (K.12), KITTI 2015 [40] (K.15), ETH3D [47] (E.), and Middlebury [46] (M.). We report D1 for KITTI, Bad 1 for ETH3D, and Bad 2 for Middlebury. We also report MACs (Multiply-Accumulate Operations) for backbone design choices. Unless otherwise specified, models are trained for 200K iterations without data augmentation on a 1.4M-image subset from synthetic datasets [61, 32, 6, 53, 38] using the default operations in [66]. Default settings in final model are marked in gray.
@@ 3.2 Training Strategy
To achieve zero-shot capability, the proposed model is trained on a large amount of high-quality data by following carefully designed training stages, which we illustrate in Fig. 4. We compile datasets into two categories: synthetic annotated data and realistic unannotated data. Since synthetic datasets provide accurate ground-truth annotations, we first train our model on them from scratch to enhance model’s matching ability. Our synthetic datasets combine SceneFlow [38] (35K), FallingThings [53] (30K), FSD [61] (1.1M), CREStereo [32] (0.2M), VKITTI2 [6] (21K), TartanAir [57] (0.31M), and Dynamic Replica [29] (0.14M), resulting in a total of 1.8M annotated samples. While other synthetic datasets are available, such as IRS [56], Sintel [5], Spring [39], and InfinigenSV [28], they are excluded due to annotation quality issues (e.g., inaccurate disparities for transparent objects in IRS) or domain mismatch (e.g., cinematic or natural scenes in Spring and InfinigenSV).
@@
where $F_{i}$ and $F^{\prime}_{i}$ are feature vectors from the teacher and student models, respectively. Here, we evaluate several distillation schemes: (a) training only the student model with fixed teacher model weights, (b) updating the teacher model via Exponential Moving Average (EMA) [43], and (c) directly copying student model weights to the teacher model at each iteration. Our ablation study (Section 4.3) shows that the simplest strategy (a) yields the best performance, and is therefore adopted in the second stage.
High-quality real-world stereo data with annotations remain scarce and sparse (e.g., Lidar-based ground truth), which constrains training scalability. In contrast, the vast amount of unlabeled real-world stereo data has not been used by the community. With the increasing availability and diversity of such data, in Stage ③, we leverage unannotated stereo pairs to improve model generalization, collecting a total of 0.5M pairs, as summarized in Tab. 1. Pseudo labels are generated using FoundationStereo [61], a strong accurate model. Importantly, even for datasets with ground-truth annotations (e.g., DrivingStereo [72]), we use pseudo but dense labels for stereo pairs (excluding the weather subset), as they provide stronger supervision than sparse data.
We further observe that data quality is more critical than scale: incorporating low-quality or domain-specific data in the third stage can degrade zero-shot performance. For example, Stereo4D [24] contains 18M stereo pairs mined from internet videos but only at a limited resolution ($\times512$), HRWSI [63] suffers from poor rectification quality, and datasets such as SCOD [41] are restricted to narrow domains. Finally, we do not apply self-distillation here, as it provides no observable performance gains when training with pseudo labels. As shown in Fig. 6, the proposed strategy progressively improves the disparity estimation, making the leaves and background regions noticeably clearer.
@@ 4 Experiments
Table 3: Zero-shot generalization results on four public benchmarks: KITTI 2012 [14], KITTI 2015 [40], ETH3D [47], and Middlebury (H) [46]. The most commonly used metrics are adopted. In the first block, all efficient methods are trained only on Scene Flow [38]. In the second and third blocks, methods are allowed to train on any existing datasets excluding the four target domains. Accuracy-based methods are shown as reference. The weights and parameters are fixed for evaluation. MACs (G) are measured at the KITTI resolution of $1242\times 375$. indicates models retrained with original code on the same synthetic data as ours. denotes results trained on 30M pseudo-labeled samples using the strategy in [19]. marks results reported by [61]. The best and second best are marked with colors.
@@
CoEX [1] 22.30 3.37 17.33 3.00 31.97 14.05 26.42 4.90 54
MobileStereoNet-2D [48] 19.30 2.51 21.88 2.85 17.10 1.83 37.98 7.54 127
MobileStereoNet-3D [48] 19.59 2.67 17.89 3.01 18.42 3.02 25.26 4.61 564
FastACV [65] 13.90 2.05 11.83 2.21 7.84 1.05 19.61 4.66 72
FastACV+ [68] 16.69 2.29 15.28 3.27 11.50 2.43 27.34 7.16 85
Lite-CREStereo++ [25] 5.93 1.29 7.37 1.36 8.95 1.18 14.91 3.32 101
LightStereo-M [20] 6.76 1.26 6.79 1.42 13.93 0.73 16.99 2.06 33
LightStereo-L [20] 6.80 1.27 6.62 1.29 9.66 0.50 17.23 2.88 84
BANet-2D [66] 14.75 2.23 16.98 3.91 44.89 35.95 26.79 6.96 36
BANet-3D [66] 17.39 2.41 17.28 3.36 29.27 14.06 28.79 8.05 78
@@
LightStereo-M [20] 4.10 0.99 4.97 1.13 5.33 0.41 10.85 1.51 33
BANet-2D [66] 3.90 0.93 4.71 1.07 5.92 0.38 10.05 1.34 36
StereoAnything-L [19] 4.00 0.92 4.81 1.10 3.81 0.31 9.82 1.21 84
@@
Selective-IGEV [58] 3.20 4.50 3.40 7.50 3619
FoundationStereo [61] 2.51 0.67 2.83 0.86 0.49 0.14 1.12 0.37 12824
Table 4: Comparison of model performance on DrivingStereo weather [72]. Lower value is better for both metrics.